Part A An internal explosion breaks an object, initially at rest, into two pieces, one of which has 2.4 times the mass of the other. If 6300 J were released in the explosion, how much kinetic energy did each piece acquire? Express your answers using two significant figures. Enter your answers numerically separated by a comma. KB Submit My Answers Give Up Incorrect, Try Again ovide Feedback Continue
Answer #1

given internal explosion in an object initially at rest
after the explosion
mass of 1st piece = m
mass of 2nd piece = 2.4 m

A. eNERGY RELEASED IN THE Explosion, E = 6300 J
   let velocity of mass m be u and that of mass 2.4m be v
   then from momentum balance
   2.4m*v = mu
   u = 2.4v

   also, from conservation of energy
   0.5mu^2 + 0.5*2.4m*v^2 = E
   0.5mu^2 + 0.5*2.4^3*m*u^2 = 6300 J
   mu^2 = 849.973

   hence KE of smaller mass KE1 = 0.5mu^2 = 424.986 J
   KE of larger mass KE2 = 0.5(2.4^3)*m*u^2 = 5875.0134 J